# Less is More : A Closer Look at Multi-Modal Few-Shot Learning

### 10 Jan 2024 v1

@@ Abstract
Few-shot Learning aims to learn and distinguish new categories with a very limited number of available images, presenting a significant challenge in the realm of deep learning. Recent researchers have sought to leverage the additional textual or linguistic information of these rare categories with a pre-trained language model to facilitate learning, thus partially alleviating the problem of insufficient supervision signals. However, the full potential of the textual information and pre-trained language model have been underestimated in the few-shot learning till now, resulting in limited performance enhancements. To address this, we propose a simple but effective framework for few-shot learning tasks, specifically designed to exploit the textual information and language model. In more detail, we explicitly exploit the zero-shot capability of the pre-trained language model with the learnable prompt. And we just add the visual feature with the textual feature for inference directly without the intricate designed fusion modules in previous works. Additionally, we apply the self-ensemble and distillation to further enhance these components. Our extensive experiments conducted across four widely used few-shot datasets demonstrate that our simple framework achieves impressive results. Particularly noteworthy is its outstanding performance in the 1-shot learning task, surpassing state-of-the-art methods by an average of 3.0% in classification accuracy.  [^1] We will make the source codes of the proposed framework publicly available upon acceptance. .
@@ 1 Introduction
Figure 1: 5-way 1-shot Performance comparisons of simple few-shot image classification baselines with the pre-trained LM. The “zero-shot” means the we train to align the visual feature and textual semantic feature, and do not use any samples from the novel classe. We utilize the trained backbone to recognize novel classes directly. The“zero-shot+LP” denotes the learnable prompts (the more details in Section 4) based on zero-shot. We observe that these two simple zero-shot baselines outperform obviously than the few-shot based ProtoNet [43] on two datasets. These experimental results demonstrate the distinguish generalization capacity of the pre-trained LM. Based on this, we proposed a simple framework leveraging the pre-trained LM directly for FSL and achieve the obvious performance improvements.
@@
Performing like humans is the ultimate goals of the Artificial Intelligent model. Recently, Deep learning-based technologies have made significant strides, achieving remarkable performances across various tasks, often rivaling or surpassing human capabilities in specific domains [26, 16, 29, 17]. However, humans have the strong ability of Few-Shot Learning (FSL) [12, 25, 49], which involves learning and discerning new classes with a very limited number of available samples. Despite the advancements in deep learning, FSL remains a significant challenge, showcasing a considerable performance gap between humans and deep learning models. At the same time, studies in Human Neuroscience provide the compelling evidence about the hypothes that humans leverage both the visual and linguistic knowledge to comprehend novel concepts and categories [20, 42, 37]. Inspired by these studies and aiming to abbreviate the problem caused by limited visual supervision signals, a series of FSL research [51, 36, 32, 6] attempt to leverage the additional modal information of available samples (e.g., textual information, also known as prompt) to assit model in recognizing new classes, which imitate the human learning processes. For instance, AM3 [51] introduces an attention based fusion mechanism to fuse the visual and textual feature, guiding the positions of class prototypes. SP-CLIP [6] proposes the semantic prompt, which utlizes the obtained textual semantic representions to guide the visual feature extraction network, employing complementary fusion mechanisms to integrate semantic representations into the feature extractor.
@@
While these works partially alleviate the problem of insufficient supervision and achieve a certain performance improvements. However, they predominantly focus on how to design intricate multi-modal fusion modules to fuse the visual and textual representations obtained by the visual encoder and textual encoder, respectively. However, these complex structures may potentially cause ignoring and even influence the generalization capacity of the used pre-trained language model (LM), leading to limited performance enhancements. At the same time, the recent research demonstrates the distinguished zero-shot capability of the pre-trained foundation language model [4, 38, 53, 34], which all trained on million-level or even billion-level language datasets. Consequently, we argue that these distinguished generalization capability of the pre-trained LM should be considered, especially for the scenarios with very limited supervision signals. To valide our motivation, we design some simple few-shot image classification baselines with the pre-trained LM, and the experimental results have been depicted in Figure 1. The zero-shot here means the we only use the base dateset to align the visual feature and textual semantic feature, without using any samples from the novel classes. Subsequently, we employ the trained backbone with pre-trained LM to recognize novel classes directly. Following previous works [38, 6], the input prompt for the LM we chose is “A photo of a [classname]”. We empirically observe that the simple zero-shot baseline outperforms obviously than the prototypical network [43] (ProtoNet in Figure 1) in the setting of 5-way 1-shot learning both on miniImageNet and CIFAR dataset, despite the fact that ProtoNet can access the extra novel samples to help to recognize novel classes. Building on this simple zero-shot baseline, we adopt the learnable prompts instead of the pre-defined fixed prompts (the details about the learnable prompts can be found in Section 4) to further improve the generalization capacity [60], denoted with “zero-shot+LP” for short. We also observe the obvious performance improvements compared to the zero-shot baseline. The detail setting of these experiments can be found in the Appendix. These experimental results valide the significant of the generalization capability of the pre-trained LM, which are usually ignored by previous FSL methods. Inspired by this insight, we aim to exploit the generalization capacity of the pre-trained LM in the FSL with very limited supervision signals.
@@
In this paper, we propose a simple and effective framework tailored for few-shot learning tasks to maximize the utilization of the textual information and pre-trained LM. We aim to utlize the generalization capacity of the pre-trained LM directly to assit the classifier for few-shot classification, instead of designing the intricate fusion modules as seen in previous works, which may ignore the generalization capacity of the obtained textual features. In more details, Our approach directly add the obtained visual feature by a visual backbone and the obtained textual semantic feature by a pre-trained LM as the Multi-modal Feature Fusion mechanism, named as SimpleFSL. Although, there are lots of recent advanced multi-modal fusion mechanism [15], we opt for the simplest Add operation in our proposed framework to valide our idea, and we argue this straightforward approach minimally impacts the generalization capability of the pre-trained LM. Notwithstanding our proposed framework has the simple network structure instead of the complex fusion mechanism, it still attains satisfactory performances. Further discussions on alternative fusion mechanisms are detailed in the experiments. Addtionally, previous works [51, 32, 6] utilize the fixed pre-defined prompts (e.g., a photo of a cat), which may constraining the generalization capability of the pre-trained LM [60]. To make a well adaptation of the pre-trained LM to diverse downstream datasets, we adopt the more flexible learnable prompts as the input of LM, automating the prompt engineering, instead of the previous fixed prompt. Concretely, We models the context words of the prompt with learnable vectors, optimized in an end-to-end way during the training period. At the same time, we still keeping pre-trained LM frozen. In this way, we enable the FSL model to learn context-aware prompts autonomously, enhancing flexibility without relying on handcrafted prompts. Furthermore, inspired by the recent advancements in Knowledge Distillation [18, 58, 57], we employ the self-ensemble and self-distillation mechanism based on the SimpleFSL to provide an additional boost for the FSL tasks, and we name this improved version as SimpleFSL++.
@@
    We explore a new perspective on multi-modal FSL, and emphasize the significance of explicitly utilization of pre-trained LM language models for FSL.
@@
    We introduce a novel and simple multi-modal few-shot learning framwork, which exploits the pre-trained language model with learnable prompts via meta-training. Further, we utilize the self-ensemble and self-Distillation to bring the additional performance improvement.
@@
    Extensive experiments across four commonly used FSL benchmarks demonstrate the satisfactory performance of the proposed simple baselines compared to state-of-the-art methods.
@@ 2.1 Few-shot Learning
In this paper, we focus on the few-shot learning task, which remains a very challenge topic in the realm of deep learning [49, 44]. The inception FSL methods such as Prototypical Network (ProtoNet) [43], Model-Agnostic Meta-Learning (MAML) [13] utilized the meta-learning strategies, aiming at acquiring transferable features. Recent advancements have focused focused on leveraging relationships among available samples to enhance discriminative representations. For instance, GNNFSL [41] pioneered the utilization of the graph nerual network [23] to explore the relations of samples. FEAT [54] established class-wise relations via the transformer [46] to derive more robust prototypical representations for inference. HGNN [55] introduced a dual-graph neural network structure to exploit the relations of samples and classes, respectively. At the same time, another line of research involves pre-training effective feature extraction networks using base datasets [5, 9, 45]. Subsequently, the pre-trained backbone will be transfered for inferring novel data. For instance, baseline++ [5] adopted the cosine similarity classifier to reduce intra-class variation among features during training. RFS [45] leveraged the Born-again strategy [14] to enhance the pre-training. SUN [10] utilized the individual supervision for local semantic learning, which helps to learn generalizable patterns in FSL.
2.2 Multi-modal Few-shot Learning
Recent endeavors in FSL have incorporated auxiliary textual information and pre-trained language models for recognizing novel classes [51, 36, 32, 6]. For instance, AM3 [51] introduced an attention based fusion mechanism to integrate the visual and textual features, guiding the learning of class prototypes. CMGNN [32] proposed a Cross-Modality Graph Neural Network to generate meta nodes with semantic information, which aids the corresponding visual feature learning. SP-CLIP [6] proposed the semantic prompts, which utlized the obtained textual semantic representions to guide the the visual feature extraction network, employing two complementary fusion mechanisms to insert semantic representions into the feature extractor. As discussed above, although these Multi-modal approaches demonstrate performance improvements, they all predominantly foucus on intricate fusion modules to leverage visual and textual information, overlooking the potential of pre-trained language models. Consequently, we explore a simple FSL framework to explicitly exploit the generalization capability of the pre-trained language model with the learnable prompt.
@@ 3.1 Problem formulation
In the context of Few-Shot Learning (FSL), a model should recognize unknown samples from unseen or novel classes with a very limited number of samples available. The dataset, which contains novel classes, is denoted as $\mathbf{D_{novel}}$. We follow the previous N-way K-shot few-shot learning setting [47, 43, 13], where a FSL task contains N classes with K labeled samples for per class, and these labeled samples are named as the support set in novel classes. The unknown or unlabeled samples in novel classes that need to be classified are denoted as the query set in each FSL task.
@@ 3.2 Meta-training
Training a FSL model directly only with very few labeled samples (i.e., one labeled sample per class) from scratch is very difficult and may cause a high risk of overfitting. Consequently, an additional base dataset with all samples annotated, denoted as $\mathbf{D_{base}}$, is provided to pre-train the FSL model to alleviate overfitting during the training stage [47, 43, 5]. In order to reduce the gap between the all labeled $\mathbf{D_{base}}$ and small partial labeled $\mathbf{D_{novel}}$, previous FSL works usually adopt the meta-training strategy, a.k.a. episodic training. In details, the meta-training strategy aims to sample a series of N-way K-shot few-shot tasks in $\mathbf{D_{base}}$. Each episode also contains a support set and a query set sampled from $\mathbf{D_{base}}$. The FSL model will be trained with $\mathbf{D_{base}}$ by the meta-training strategy until the FSL model converges. Note that the classes in $\mathbf{D_{base}}$ and $\mathbf{D_{novel}}$ are entirely disjoint, and there is no any overlap in the class label spaces or samples between them. Thus the FSL model should have the ability to generalize to unseen classes with few labeled samples.
@@ 4 Method
Figure 2: The summary of our proposed SimpleFSL and SimpleFSL++. Instead of designing the intricate fusion modules as seen in previous works, we directly add the obtained visual feature by a visual backbone and the obtained textual semantic feature for few-shot classification tasks. We also adopt the more flexible learnable prompts as the input of LM to automate prompt engineering, instead of the previous fixed prompt. We name this method as SimpleFSL, and we further utilize the self-Ensemble and Distillation to improve performance (the dotted box part), named as SimpleFSL++
In this section, we introduce the details of our proposed framwork as illustrated in Figure 2. The whole framwork is very simple, containing three modules: visual backbone for visual feature extraction from the images, textual backbone with learnable prompts for textual semantic feature extraction, and feature fusion module for the final prediction. The SimpleFSL++ contains the additional self-ensemble and self-distillation module to improve the performance further.
@@ 4.1 Pre-training
Before the meta-training, we first pre-train our visual backbone with $\mathbf{D_{base}}$ to accelerate the convergence, following previous work [54, 7, 6]. During the pre-training stage, the visual features of samples in $\mathbf{D_{base}}$ will be obtained by a visual backbone $f_{\Theta}(\ )$ (e.g., Visual Transformer [11]) with learnable parameters $\mathbf{\Theta}$. After that, we use a simple linear classifier with learnable parameters $\mathbf{\{W,b\}}$ containing a weight term $\mathbf{W}$ and a bias term $\mathbf{b}$, which maps the input feature into one of the base classes. These processes are optimized by minimizing the standard cross entropy loss, and formally we have:
$\mathcal{L}_{\text{pre }}=\frac{1}{\left|D_{\text{base }}\right|}\sum_{(\boldsymbol{x},y)\in D_{\text{base }}}-\log\frac{\exp\left(\boldsymbol{W}_{y}^{T}f(\boldsymbol{x})+\boldsymbol{b}_{y}\right)}{\sum_{i}\exp\left(\boldsymbol{W}_{i}^{T}f(\boldsymbol{x})+\boldsymbol{b}_{i}\right)}+\mathcal{R}$ (1)
where $\boldsymbol{x}$ denotes the image in $\mathbf{D_{base}}$ and $y$ denotes the ground-truth label. And $\mathcal{R}$ is the L-2 regularization term.

### 24 Mar 2024 v2

# Less is More: A Closer Look at Semantic-based Few-Shot Learning

@@ Abstract
Few-shot Learning aims to learn and distinguish new categories from a scant number of images, presenting a significant challenge in the realm of deep learning. Recent researchers have sought to leverage the additional semantic or linguistic information of scarce categories with a pre-trained language model to facilitate learning, thus partially alleviating the problem of insufficient supervision signals. Nonetheless, the full potential of the semantic information and pre-trained language model have been underestimated in the few-shot learning till now, resulting in limited performance enhancements. To address this, we propose a straightforward and efficacious framework for few-shot learning tasks, specifically designed to exploit the semantic information and language model. More specifically, we explicitly harness the zero-shot capability of the pre-trained language model with learnable prompts. And we directly add the visual feature with the textual feature for inference without the intricate designed fusion modules as in prior studies. Additionally, we apply the self-ensemble and distillation to further enhance performance. Extensive experiments conducted across four widely used few-shot datasets demonstrate that our simple framework achieves impressive results. Particularly noteworthy is its outstanding performance in the 1-shot learning task, surpassing the current state-of-the-art by an average of 3.3% in classification accuracy.  [^1] We will make the source codes of the proposed framework publicly available upon acceptance.
Keywords:
Few-Shot Learning Image Classification
@@ 1 Introduction
Performing like humans is the ultimate goals of the Artificial Intelligent models. Recently, Deep learning-based technologies have made significant strides, achieving remarkable performances across various tasks, often rivaling or surpassing human capabilities in specific domains [28, 18, 31, 17]. However, humans have the strong ability of Few-Shot Learning (FSL) [12, 26, 56], which involves learning and discerning new classes with a very limited number of available samples. Despite the advancements in deep learning, FSL remains a significant challenge, showcasing a considerable performance gap between humans and deep learning models. At the same time, studies in Human Neuroscience provide the compelling evidence about the hypothes that humans leverage both the visual and linguistic knowledge to comprehend novel concepts and categories [21, 48, 42]. Inspired by these studies and aiming to alleviate the problem caused by limited visual supervision signals, a series of FSL research [58, 40, 34, 6] attempt to leverage the additional Semantic information of available samples (e.g., textual information, also known as prompt) to assist models in recognizing rare classes, which imitate the human learning processes. For instance, AM3 [58] introduces an attention based fusion mechanism to fuse the visual and textual feature, guiding the positions of visual class prototypes. SP-CLIP [6] proposes the semantic prompt, which utilizes the obtained textual semantic representations to guide the visual feature extraction network, employing two sophisticated complementary fusion mechanisms to integrate semantic representations into the feature extractor.
Figure 1: 5-way 1-shot Performance comparisons of simple few-shot image classification baselines with the pre-trained LM. The "zero-shot" align the visual feature and textual semantic feature, and do not use any samples from the novel classes. We utilize the trained backbone to directly recognize novel classes. The "zero-shot+LP" denotes the learnable prompts (more details in Section 4) based on zero-shot. Two simple baselines outperform obviously than ProtoNet [49], showing the distinguish generalization capacity of LMs in FSL.
@@
While these works [58, 40, 34, 60, 6] partially alleviate the problem of insufficient supervision and achieve a certain performance improvements, they predominantly focus on how to design intricate multi-modal fusion modules to fuse the visual and semantic representations obtained by the visual encoder and textual encoder, respectively. However, these complex structures may potentially cause ignoring and even influence the generalization capacity of the used pre-trained language model (LM), leading to limited performance enhancements. At the same time, the recent research demonstrates the distinguished zero-shot capability of the pre-trained foundation language model [4, 44, 61, 38], which all trained on million-level or even billion-level language datasets. Consequently, we argue that these distinguished generalization capability of the pre-trained LM should be considered, especially for the scenarios with very limited supervision signals. To validate our motivation preliminarily, we design two simple few-shot image classification baselines with the pre-trained LM, and the experimental results have been depicted in Figure 1. The "zero-shot" here means the we only use the base dateset to align the visual feature and textual semantic feature, without using any samples from the novel classes. Subsequently, we employ the trained backbone with pre-trained LM to recognize novel classes directly without training. Following previous works [44, 6], the input prompt for the LM we chose is "A photo of a [classname]". We empirically observe that the simple zero-shot baseline outperforms obviously than the prototypical network [49] (ProtoNet in Figure 1) in the setting of 5-way 1-shot learning both on miniImageNet and CIFAR dataset, despite the fact that ProtoNet can access the extra novel samples to help to recognize novel classes. Further, building on this simple zero-shot baseline, we adopt the learnable prompts instead of the pre-defined fixed prompts (the details about the learnable prompts can be found in Section 4) to further improve the generalization capacity [68], denoted with "zero-shot+LP" for short. We also observe the obvious performance improvements compared to the zero-shot baseline. The detail setting of these experiments can be found in the Appendix. These experimental results validate our motivation preliminarily and show the significant of the generalization capability of the pre-trained LM, which are usually ignored by previous FSL methods. Inspired by this phenomenon, we aim to exploit the generalization capacity of the pre-trained LM in the FSL with very limited supervision signals.
@@
Consequently, we propose a straightforward and efficacious framework tailored for few-shot learning tasks to maximize the utilization of the textual information and pre-trained LM in this paper. We aim to utilize the generalization capacity of the pre-trained LM directly to assist the classifier for few-shot classification, instead of designing the intricate and complex fusion modules as seen in previous works, which may ignore and even hurt the generalization capacity of the semantic features obtained by LMs. In more details, our approach directly adds the visual feature obtained by a visual backbone and the textual semantic feature obtained by a pre-trained LM as the Multi-modal Feature Fusion mechanism, named as SimpleFSL. Although, there are lots of recent advanced multi-modal fusion mechanism [15], we opt for the simplest Add operation in our proposed framework to validate our idea, and we argue this straightforward approach minimally impacts the generalization capability of the pre-trained LM. Despite having a simple network structure instead of the complex fusion mechanism, our proposed framework still attains satisfactory performances. Additional discussions on alternative fusion mechanisms are detailed in the experiments. Addtionally, previous works [58, 34, 6] utilize the fixed pre-defined prompts (e.g., a photo of a cat), which may constraining the generalization capability of the pre-trained LM [68]. To make a well adaptation of the pre-trained LM to diverse downstream datasets, we adopt the more flexible learnable prompts as the input of LM, automating the prompt engineering, instead of the previous fixed prompt. Concretely, We model the context words of the prompt with learnable vectors, optimized in an end-to-end way via meta-training [13]. At the same time, we still keeping pre-trained LM frozen. In this way, we enable the FSL model to learn context-aware prompts autonomously, enhancing the flexible accommodation to various datasets without relying on handcrafted prompts. Furthermore, inspired by the recent advancements in Knowledge Distillation [19, 66, 65], we employ the self-ensemble and self-distillation mechanism based on the SimpleFSL to provide an additional boost for the FSL tasks, and we name this improved version as SimpleFSL++.
@@
    We explore a new perspective on Semantic-based FSL, and emphasize the significance of explicitly utilization of pre-trained language models for FSL.
@@
    We introduce a novel and simple semantic-based few-shot learning framework, which exploits the pre-trained language model with learnable prompts via meta-learning. Further, we utilize the self-ensemble and self-Distillation to bring the additional performance improvement.
@@
    Extensive experiments across four commonly used FSL benchmarks demonstrate the satisfactory performance of the proposed simple baselines compared to the state-of-the-art methods.
@@ 2.1 Few-shot Learning
In this paper, we focus on the few-shot learning task, which remains a very challenge topic in the realm of deep learning [56, 50]. The inception FSL methods such as Prototypical Network (ProtoNet) [49], Model-Agnostic Meta-Learning (MAML) [13] utilize the meta-learning strategies, aiming at acquiring transferable features. Further, Meta-AdaM [51] proposes a mata-learned learning rate learner for more rapid convergence compared to MAML. Some recent advancements have pivoted towards leveraging relationships among available samples to enhance discriminative feature representation. For instance, GNNFSL [47] pioneers the utilization of the graph nerual network [24] to explore the relations of samples. FEAT [62] establishes class-wise relations via transformers [53] to derive more robust prototypical representations for inference. HGNN [63] introduces a dual-graph neural network structure to exploit the relations of samples and classes, respectively. Concurrently, another line of research involves pre-training effective feature extraction networks on base datasets [5, 9, 52], which then are transferred for inferring novel data. For instance, baseline++ [5] adopted the cosine similarity classifier to reduce intra-class variation among features during training. RFS [52] leverages the Born-again strategy [14] to enhance pre-training. SUN [10] utilizes the individual supervision for local semantic learning, which helps to learn generalizable patterns in FSL.
2.2 Semantic-based Few-shot Learning
Recent endeavors in FSL have incorporated auxiliary semantic information and pre-trained language models to assit the recognition of novel classes [58, 40, 34, 60, 6]. For instance, AM3 [58] introduces an attention based fusion mechanism to integrate the visual and textual features, guiding the learning of class prototypes. CMGNN [34] proposes a Cross-Modality Graph Neural Network to generate meta nodes with semantic information, which aids the corresponding visual feature learning. SP-CLIP [6] proposed the semantic prompts, which utlizes the obtained textual semantic representations to guide the the visual feature extraction network, employing two complementary fusion mechanisms to insert semantic representations into the feature extractor. As discussed above, although these semantic-based approaches have shown performance gains, they all predominantly concentrate on intricate fusion modules to leverage visual and textual information, inadvertently overlooking the potential of pre-trained language models. Consequently, we explore a straightforward FSL framework to explicitly exploit the generalization capability of the pre-trained language model with learnable prompts.
@@ 3.1 Problem formulation
In the context of Few-Shot Learning (FSL), the objective of a model is to recognize unknown samples from unseen or novel classes via leveraging a very limited number of available samples. Formally, the dataset encompassing novel classes, is denoted as $\mathbf{D_{novel}}$. We follow the previous N-way K-shot learning setting [54, 49, 13], where an FSL task comprises N classes with K labeled samples for per class, and these labeled samples are named as the support set in novel classes. Conversely, the unknown or unlabeled samples in novel classes, which the model aims to classify, are denoted as the query set in each FSL task.
@@ 3.2 Meta-training
Training an FSL model directly from scratch only with the limited number of labeled samples (i.e., one per class) poses significant difficulties and may cause a high risk of overfitting. Consequently, an additional base dataset with all samples annotated, denoted as $\mathbf{D_{base}}$, is provided to pre-train the FSL model to alleviate overfitting in the training phase [54, 49, 5]. In order to reduce the gap between the full labeled $\mathbf{D_{base}}$ and very few partial labeled $\mathbf{D_{novel}}$, prevalent FSL works usually adopt the meta-training strategy [13], a.k.a. episodic training. In details, meta-training endeavors to sample a series of N-way K-shot learning tasks per episode from $\mathbf{D_{base}}$. Each episode also contains a support set and a query set sampled from $\mathbf{D_{base}}$, as have been done from $\mathbf{D_{novel}}$. The FSL model will be trained with $\mathbf{D_{base}}$ via meta-training until the FSL model converges. It is crucial to note that the class categories in $\mathbf{D_{base}}$ and $\mathbf{D_{novel}}$ are entirely disjoint, without any overlap in the their label spaces or samples, thus ensuring the model’s capacity for generalizing to unseen classes with limited labeled data.
@@ 4 Method
Figure 2: The schematic of our proposed SimpleFSL and SimpleFSL++. Instead of designing the intricate fusion modules as seen in previous works, we directly add the obtained visual feature by a visual backbone with the obtained textual semantic feature for few-shot classification tasks. We also adopt the more flexible learnable prompts as input to the LM to automate prompt engineering, instead of the previous fixed prompts. We name this method as SimpleFSL, and we further utilize the self-ensemble and self-distillation to further improve performance (the dotted box part), encapsulated in SimpleFSL++.
In this section, we introduce the details of our proposed framework SimpleFSL and its variants SimpleFSL++, as both depicted in Fig.2. The whole framework is very simple without complex architectures, which contains three primary modules: a visual backbone for visual feature extraction from the images, a textual backbone with input prompts for the textual semantic feature extraction, and feature fusion module for the final prediction. Additionally, SimpleFSL++ contains the self-ensemble and self-distillation module to improve the performance further.
@@ 4.1 Pre-training
Following previous works [62, 7, 6], we first pre-train our visual backbone on $\mathbf{D_{base}}$ to expedite convergence, before the meta-training. During the pre-training stage, the visual features of all samples in $\mathbf{D_{base}}$ are extracted by a visual backbone $f_{\Theta}(\ )$ (e.g., Visual Transformer [11]) with learnable parameters $\Theta$. Subsequently, we use a simple linear classifier with learnable parameters $\mathbf{\{W,b\}}$ comprising a weight term $\mathbf{W}$ and a bias term $\mathbf{b}$, which maps the input features into one of the base classes. This procedure is optimized by minimizing the standard cross entropy loss, described formally as follows:
$\small\mathcal{L}_{\text{pre }}=\frac{1}{\left|D_{\text{base }}\right|}\sum_{(\boldsymbol{x},y)\in D_{\text{base }}}-\log\frac{\exp\left(\boldsymbol{W}_{y}^{T}f(\boldsymbol{x})+\boldsymbol{b}_{y}\right)}{\sum_{i}\exp\left(\boldsymbol{W}_{i}^{T}f(\boldsymbol{x})+\boldsymbol{b}_{i}\right)}+\mathcal{R}$ (1)
where $\boldsymbol{x}$ represents an image in $\mathbf{D_{base}}$, $y$ is the corresponding ground-truth label, and $\mathcal{R}$ denotes the L-2 regularization term.